# Reducing Size Bias in Sampling for Infectious Disease Spread on Networks

### v1

@@ Abstract
Epidemiological models can aid policymakers in reducing disease spread by predicting outcomes based on disease dynamics and contact network characteristics. Accurate calibration of these models requires representative samples of contact networks. In this connection, we investigate the performance of two sampling algorithms—Random Walk (RW) and Metropolis-Hastings Random Walk (MHRW)—across three network types: Erdős–Rényi (ER), Small-World (SW), and Scale-Free (SF). Disease transmission is simulated using a stochastic susceptible-infected-recovered (SIR) framework. Our findings show that for ER and SW networks, RW overestimates infected individuals and secondary infections by approximately $25\%$, while MHRW produces estimates that align well with the underlying network structure. Both methods also yield estimates closer to the underlying network for time-to-infection. However, sampling SF networks introduces significant variability for both algorithms. Removing duplicate samples further reduces MHRW’s accuracy in all network types. We apply both algorithms to a cattle movement network consisting of 46,512 farms, exhibiting characteristics of ER, SW, and SF networks. RW overestimates infected farms by approximately $100\%$ and secondary infections by $>900\%$, whereas MHRW reduces these overestimations to approximately $50\%$ and $<800\%$, respectively. Time-to-infection estimates reveal that RW underestimates by approximately $40\%$, while MHRW slightly overestimates by $10\%$. These results highlight the importance of selecting sampling algorithms based on network structure and disease severity. RW provides conservative estimates suited for high-mortality, fast-spreading diseases, while MHRW is more suitable for precise interventions in less severe outbreaks. These insights can guide policymakers in optimising resource allocation and disease control strategies.
@@
[inst1]organization=School of Mathematics, Cardiff Univerity,addressline=Senghennydd Road, city=Cardiff, postcode=CF24 4AG, country=UK
@@ 1 Introduction
During an epidemic, accurately estimating metrics such as the number of infected individuals, effective reproduction number, and time to infection is crucial for preparedness (Nunes et al., 2024). Mathematical modelling provides a suite of tools, frameworks, and methodologies to generate these estimations (Vynnycky and White, 2010). However, estimating disease metrics on networks is challenging due to factors such as network complexity (Avraam and Hadjichrysanthou, 2025) and the influence of the chosen data sampling method (Danon et al., 2011). In particular, the Random Walk (RW) sampling algorithm is prone to size bias, leading to over-representation of highly connected individuals and potentially skewed estimates of disease spread. In this paper, we statistically compare two sampling algorithms, Random Walk (RW) and Metropolis-Hastings Random Walk (MHRW), using stochastic susceptible-infected-recovered (SIR) model simulations to evaluate the estimation of disease metrics on three network types (Erdős-Rényi (ER), Small-world (SW), and Scale-free (SF)) and discuss the implications for policy planning.
@@
We consider three types of contact networks: 1) ER networks (Erdős and Rényi, 1959), which serve as a simple yet useful benchmark for understanding disease spread in random, unstructured networks; 2) SW networks (Watts and Strogatz, 1998), which closely mimic social mixing networks due to the high clustering and long-range links; and 3) SF networks (Li et al., 2005), characterized by the presence of hubs with high connectivity, reflecting the presence of super-spreaders in a real-world contact network (Lieberthal and Gardner, 2021). SF networks capture the extreme heterogeneity of connections found in real-world contact networks, which is less pronounced in the SW and ER networks, which have relatively uniform connectivity. Although these are theoretical networks they capture many properties of real-world networks (Aldous and Wilson, 2003). Thus, we use them here to demonstrate the diversity of behaviours that can arise.
@@
Since our primary interest is to understand the influence of the underlying contact network on disease transmission metrics, we consider a stochastic SIR model (Brauer, 2008) of disease spread on the contact network. The SIR models are the most commonly used mechanistic compartmental models for infectious disease spread (Kermack and McKendrick, 1927). SIR dynamics can be simulated at either the population level or the individual level; here we do the latter.
@@
A SIR model at the population level consists of a deterministic system of ordinary differential equations (ODEs), which describes the rate of change in the number or proportion of individuals in three compartments: susceptible, infected, and recovered, at the population level. The solution of ODEs predict key features of disease spread, such as the number of infections, the rate of transmission, and the effective reproduction number (Brauer, 2008).
@@
The SIR dynamics at the individual level, as we use in this work, takes the form of stochastic equations describing the rate of change in the probability of the population’s state. The SIR framework at the individual level, accounts for heterogeneous mixing within the population based on the underlying contact network (Kiss et al., 2017a).
@@
Another approach to studying disease spread is to use data to extract correlation patterns using statistical modelling. In statistical models, usually, only a sample of the data is adequate. There are two key reasons for using sample data: a) the large cost and time needed in collecting data for an entire population, and b) the high computational cost of processing large datasets. The accuracy of predictions of both statistical and mechanistic models depends on the quality of the sample; if the sample is not representative of the population then it leads to biased estimates of disease metrics.
@@
Various types of biases can arise if an incorrect sampling method is employed, including demographic bias (Tyrer and Heyman, 2016), geographic bias (Banerjee and Chaudhury, 2010), size bias, and temporal bias (Kleinbaum et al., 1981). For instance, in cohort studies for survival analysis of cancer patients, size bias occurs when individuals with longer survival times are more likely to be included in the sample. This can lead to an overestimation of survival time if the bias is not corrected (Shen et al., 2009).
@@
For diseases which are highly transmissible via close contact such as COVID-19, measles, chickenpox etc., the number of contacts is a critical factor in the spread of the virus, it is essential to obtain an unbiased sample that accurately reflects the population’s contact distribution. Failure to do so can lead to incorrect estimation and prediction of the disease spread rate and ineffective mitigation.
@@
Two classic probabilistic TBS methods are the RW (Wei et al., 2004) and MHRW algorithms (Hu and Lau, 2013). The RW and MHRW are Markov chain-based algorithms and share the “memoryless” property, leading to duplicates in the sample. In this work, we also compare the sampling algorithms when we remove or keep the duplicate sample points.
The RW algorithm is often used in epidemiological studies of large populations (Milligan et al., 2004, Qi et al., 2023) as it is relatively fast compared to other sampling algorithms. Specifically, for infectious diseases, the RW algorithm works well for homogeneous contact networks, where individuals have a similar number of contacts. However, real-world contact networks are often heterogeneous (Nielsen et al., 2020), that is, the individuals vary in their number of contacts. This heterogeneity can result in size biased samples when using the RW algorithm, i.e., over-representation of highly connected individuals (high-degree nodes) (Stein et al., 2014, Li et al., 2015).
In (Gjoka et al., 2010), the authors compare the RW and MHRW algorithms for sampling online social networks, specifically Facebook. Their results show that the RW algorithm produces a size biased sample, over-representing high-degree nodes, whereas the MHRW algorithm generates a sample with a degree distribution closely matching that of the underlying Facebook network. Similar observations are reported in (Leskovec and Faloutsos, 2006b) for various network types, including citation networks from the e-print arXiv (Leskovec and Faloutsos, 2006b), autonomous system networks (University of Oregon, ) (Internet router networks), affiliation networks from the astrophysics category on arXiv (Leskovec and Faloutsos, 2006b), and trust networks from Epinions.com (Richardson et al., 2003). We note, however, that similar comparative analyses in the context of epidemiological research are limited. This study seeks to address this shortcoming by exploring the RW and MHRW sampling methods in the domain of infectious disease modelling.
Here, we examine how network structures and sampling algorithms affect estimation of disease metrics. We, thus, simulate a stochastic SIR model with varying transmission and recovery rates for three network types (ER, SW and SF). Specifically, in Section 2, we illustrate the size bias in the measurement of disease metrics introduced by the RW and MHRW algorithms, and in Section 3 we compare key metrics such as epidemic size, the number of secondary infections (effective reproduction number), and the time to infection. Finally, in Section 4, to illustrate the practical implications of our study, we analyse cattle movement network data provided by the British Cattle Movement Service (BCMS), which has hybrid properties linked to the ER, SW and SF networks. We provide a summary and directions for future research in Section 5.
@@ 3.2 Disease metrics
In the previous section, we compared the degree distribution of the samples generated using the RW and MHRW algorithms; we observed that the degree distribution of MHRW samples is closer to the UN than RW samples, due to bias towards high-degree nodes. Now, we compare the sampling algorithms in estimating three disease metrics: 1) proportion of infected nodes, 2) average number of secondary infections, and 3) time to infection. We simulate the SIR model on networks to estimate the disease metrics in the samples and the UN. Here, we present the results for $\gamma=1$; see C.2 for $\gamma\in\{1/7,1/14\}$.
@@ 3.2.2 Average secondary infections
Results for the time to infection disease metric are described in C.1. We observed that in Figure 8, estimates from both sampling algorithms (with or without duplicates) are close to the UN. Although the RW samples slightly underestimate the time to infection, whereas the MHRW samples align well with the UN.
@@
The RW samples overestimated the disease metrics regardless of whether duplicate nodes are removed or retained. The size bias in degree distribution (Section 3.1), for the RW samples, leads to overestimating the number of secondary infections for all networks, as the nodes with high degrees are more likely to become infected.
@@
From an application perspective, the RW sampling algorithm provides conservative estimates of disease metrics, which is beneficial for devising interventions in cases of infectious diseases with high fatality rates. For such diseases, e.g. COVID-19 and Ebola, it is preferable to contain transmission as early as possible; hence, overestimating the number of infected individuals and the effective reproduction number (i.e., the number of secondary infections) aids the implementation of timely prevention measures. However, for diseases with low fatality rates, policymakers have more time to develop policies aligned with available resources. Here, the MHRW sampling algorithm is preferable, as its estimates are closer to the UN structure. In this context, precise estimates of disease metrics is integral for efficient use of time and resources to contain the disease spread.
@@ 4 Cattle Movement Network
In Section 3.2, we compared the network properties and disease metrics estimated from the RW and the MHRW sampling algorithms, on three theoretical network types. Now, we implement both sampling algorithms on movement network data. We use the cattle movement data provided by the British Cattle Movement Service (BCMS), which contains information about birth, death, and movement of the cattle.
@@ 4.1 Data Description
The data contains movement details of $1,135,502$ cattle; there are two movement records on average, per cattle. There are $46,512$ holding places, which are the nodes of the network. A static and undirected network is created with $46,512$ nodes and $159,036$ edges. Note that an edge between a pair of nodes is counted only once, in case of multiple movement records. A part of the movement network is shown in Figure 6(a); where we see a mix of large connected component, small clusters of nodes and isolated node pairs.
@@
(e) Time to infection (days)
@@
(h) Time to infection (days)
@@ 4.2 Results
In Figure 7, we compare three disease metrics in RW and MHRW samples relative to the cattle network, as the transmission rate $\beta$ increases, with a recovery rate $\gamma=1$, for the cattle movement network. The three disease metrics are: 1) proportion of infected nodes, 2) number of secondary infections, and 3) time to infection. Panel 1 (Figure 6(c), 6(d), 6(e)) shows the results for samples with duplicate nodes, while Panel 2 (Figure 6(f), 6(g), 6(h)) shows those without duplicates. The closer the disease metric estimates made from the samples are to the solid horizontal black line, the closer they are to the underlying network.
@@
In Panel 1 of Figure 7, the estimate of time to infection is overestimated by the MHRW samples and underestimated by the RW samples. The variability in estimates is extremely high for the MHRW samples compared to the thin band for the RW sample estimates (Figure 6(e)). Interestingly, after removing the duplicate nodes, both sampling algorithms underestimate the time to get the infection, which deviates further from the cattle network estimate as the $\beta$ value increases.
Figure 7 indicates that the estimates from MHRW samples are closer to the cattle network than the RW samples. However, the variability is significantly higher for the MHRW samples for the number of secondary infections and the time to infection.
@@ 5 Conclusions
In this study, we examined the performance of the Random Walk (RW) and Metropolis-Hastings Random Walk (MHRW) sampling algorithms for estimating disease metrics in three network types—Erdős-Rényi (ER), Small-world (SW), and Scale-free (SF). Using a stochastic SIR model, we investigated the effects of sampling biases, particularly size bias, and the inclusion or exclusion of duplicate nodes on estimates of disease spread metrics. Our analysis focused on three disease metrics: the proportion of infected individuals, the average number of secondary infections, and the time to infection.
@@
Our findings (Section 3.2) highlight the suitability of the MHRW algorithm for networks with properties similar to the ER and SW networks. MHRW samples consistently produced estimates that closely aligned with the underlying network structure, particularly when duplicate samples were retained. In contrast, the estimates from the RW samples exhibited a significant overestimation of the proportion of infected individuals and secondary infections, while underestimating the time to infection. This makes the RW algorithm less reliable for precise estimates but useful for generating conservative predictions, which can be valuable for early interventions in high-fatality diseases like COVID-19 and Ebola.
@@
In SF networks, the inherent heterogeneity and presence of high-degree hubs introduced substantial variability in the estimates derived from both the RW and MHRW samples. This suggests that neither method is optimal for such networks. Alternative approaches, such as separate sampling of core and peripheral nodes using algorithms like SLSR (Jiao, 2024), may be required for more accurate and balanced sampling.
@@
The cattle movement network, which exhibits characteristics of ER, SW, and SF networks, further supported these observations. MHRW samples were more representative of the underlying network when duplicate nodes were retained, while RW samples tended to overestimate or underestimate disease metrics. For highly contagious diseases like Foot-and-Mouth Disease, the conservative estimates from RW sampling are advantageous for rapid containment measures. For less severe diseases like Bovine Viral Diarrhea, the more precise estimates from MHRW sampling better support resource-efficient intervention strategies.
@@
Additionally, our analysis revealed that the sensitivity of both algorithms to recovery rates ($\gamma$) significantly impacts disease metric estimates (C.2). While deviations in the proportion of infected nodes diminish with lower recovery rates, variability in the estimates of secondary infections and time to infection increases with higher transmission rates ($\beta$). These findings emphasise the importance of accounting for disease-specific parameters when choosing a sampling algorithm.
@@
To further understand the subtle differences between the RW and MHRW sampling algorithms with respect to network structure, it would be valuable to explore the impact of selecting a specific starting node for the sampling process, rather than a random selection, particularly for SF networks. As observed, the difference in disease estimates between the RW and MHRW samples decreases with a decrease in the recovery rate. It would be interesting to determine the value of the recovery rate at which this change starts, as it could guide the choice of sampling algorithm. Additionally, exploring the sensitivity of these estimates to variations in network generation parameters, such as network size, average degree, and the power-law exponent for SF networks, would provide further insight.
@@
Conceptualization, Formal analysis, Investigation, Methodology, Validation, Visualization, Writing – original draft & editing. Katerina Kaouri: Conceptualization, Formal analysis, Investigation, Methodology, Project administration, Supervision, Validation, Writing – review & editing. Thomas Woolley: Conceptualization, Formal analysis, Investigation, Methodology, Project administration, Supervision, Validation, Writing – review & editing.

### v2

# Reducing Size Bias in Sampling Infectious Disease Spread on Networks

@@ Abstract
Epidemiological models can aid policymakers in reducing disease spread by predicting outcomes based on disease dynamics and contact network characteristics. Calibrating these models requires representative network samples. In this connection, we investigate two sampling algorithms, Random Walk (RW) and Metropolis-Hastings Random Walk (MHRW), across three network types: Erdős-Rényi (ER), Small-World (SW), and Scale-Free (SF). Disease transmission is simulated using a stochastic susceptible-infected-recovered (SIR) framework. Our findings show that RW overestimates infected individuals and secondary infections by $25\%$ for ER and SW networks due to size bias, favouring highly connected nodes. MHRW, which corrects for size bias, provides estimates that are more consistent with the underlying network. Also, both methods yield estimates closer to the underlying network for time-to-infection. However, sampling SF networks exhibits significant variability, for both algorithms. Removing duplicate sampled nodes reduces MHRW’s accuracy across all network types. We apply both algorithms to a cattle movement network of 46,512 farms, exhibiting ER, SW, and SF network features. RW overestimates infected farms by approximately $100\%$ and secondary infections by $>900\%$, reflecting size bias whereas MHRW estimates align closely with the cattle network dynamics. Time-to-infection estimates reveal that RW underestimates by approximately $40\%$, while MHRW slightly overestimates by $10\%$. Estimates differ greatly when duplicate nodes are removed. These findings underscore choosing algorithms based on network structure and disease severity. RW’s conservative estimates suit high-mortality, fast-spreading diseases, while MHRW provides precise interventions suitable for less severe outbreaks. These insights can guide policymakers in optimising resource allocation and disease control strategies.
@@
[inst1]organization=School of Mathematics, Cardiff University,addressline=Senghennydd Road, city=Cardiff, postcode=CF24 4AG, country=UK
{highlights}
Assessed RW and MHRW sampling for disease metrics in ER, SW, and SF networks.
MHRW outperformed RW in ER and SW, reducing size bias for accurate estimates.
SF networks showed high variability, suggesting alternative sampling methods.
RW sampling overestimated disease metrics, while MHRW yielded precise results.
@@ 1 Introduction
During an epidemic, accurately estimating metrics such as the number of infected individuals, effective reproduction number, and time-to-infection is crucial for effective mitigation (Nunes et al., 2024). Mathematical modelling provides a suite of tools, frameworks, and methodologies to generate these estimations (Vynnycky and White, 2010). However, estimating disease metrics on networks is challenging due to factors such as network complexity (Avraam and Hadjichrysanthou, 2025) and the influence of the chosen data sampling method (Danon et al., 2011). The Random Walk (RW) algorithm, a commonly used sampling method, is prone to size bias, leading to over-representation of highly connected nodes and potentially skewed estimates of disease spread. Metropolis-Hastings Random Walk (MHRW) addresses this limitation by correcting for size bias, offering a more balanced representation of network nodes.
In this paper, we statistically compare RW and MHRW to evaluate their performance in estimating disease metrics. Using stochastic susceptible-infected-recovered (SIR) simulations, we assess these algorithms across three network types, Erdős-Rényi (ER), Small-World (SW), and Scale-Free (SF), to identify their strengths and limitations. Additionally, we discuss the implications of our findings for policy planning, highlighting how sampling methods influence the accuracy of disease spread estimates and, consequently, the effectiveness of mitigation strategies.
@@
We consider three types of contact networks: 1) ER networks (Erdős and Rényi, 1959), which serve as a simple yet useful benchmark for understanding disease spread in random, unstructured networks like sexual contact networks (Holme, 2013); 2) SW networks (Watts and Strogatz, 1998), which closely mimic social mixing networks due to the high clustering and long-range links; and 3) SF networks (Li et al., 2005), characterized by the presence of hubs with high connectivity, reflecting the presence of super-spreaders in a real-world contact network (Lieberthal and Gardner, 2021). SF networks capture the extreme heterogeneity of connections found in real-world contact networks, which is less pronounced in the SW and ER networks, that have relatively uniform connectivity. Although these are theoretical networks they capture many properties of real-world networks (Aldous and Wilson, 2003). Thus, we use them here to demonstrate the diversity of behaviours that can arise.
@@
Since our primary interest is to understand the influence of the underlying contact network on disease transmission metrics, we consider a stochastic SIR model (Brauer, 2008) to simulate disease spread on the contact network. SIR models are the most commonly used mechanistic compartmental models for infectious disease spread (Kermack and McKendrick, 1927). SIR dynamics can be simulated at either the population level or the individual level; here we do the latter.
@@
A SIR model at the population level consists of a deterministic system of ordinary differential equations (ODEs), which describes the rate of change in the number or proportion of individuals in three compartments: susceptible, infected, and recovered. The solution predicts key features of disease spread, such as the number of infections, the rate of transmission, and the effective reproduction number (Brauer, 2008).
@@
The SIR dynamics at the individual level, employed here, take the form of stochastic equations describing the rate of change in the probability of a population in a certain state and can account for heterogeneous mixing within the population based on the underlying contact network (Kiss et al., 2017a).
@@
Another approach to studying disease spread is to use statistical modelling for extracting correlation patterns from data. In statistical models, usually, only a sample of the data is required. There are two key reasons for using sample data: a) the large cost and time needed in collecting data for an entire population, and b) the high computational cost of processing large datasets. The accuracy of predictions of both statistical and mechanistic models depends on the quality of the sample; if the sample is not representative of the population then it leads to biased estimates of disease metrics.
@@
Various biases can arise from using inappropriate sampling methods, including demographic bias (Tyrer and Heyman, 2016), geographic bias (Banerjee and Chaudhury, 2010), temporal bias (Kleinbaum et al., 1981), and size bias (Arratia et al., 2019). In this study, we focus on size bias, which occurs when nodes with higher connectivity are sampled disproportionately often. For example, in survival analysis cohort studies of cancer patients, size bias can result in individuals with longer survival times being overrepresented. If uncorrected, this skews results, leading to an overestimation of survival times (Shen et al., 2009).
@@
For diseases which are highly transmissible via close contact, such as COVID-19, measles, chickenpox etc., the number of contacts is a critical factor in the spread of the virus and it is essential to obtain an unbiased sample that accurately reflects the population’s contact distribution. Failure to do so can lead to size bias prediction of the disease spread rate and consequently ineffective mitigation.
@@
Two classic and frequently used probabilistic TBS methods are the RW (Wei et al., 2004) and MHRW algorithms (Hu and Lau, 2013). The RW and MHRW, as Markov chain-based algorithms share the “memoryless” property, leading to duplicates in the sample. In this work, we also compare these sampling algorithms when we remove or keep duplicate sample points.
The RW algorithm is often used in epidemiological studies of large populations (Milligan et al., 2004, Qi et al., 2023) as it is relatively cheap compared to other sampling algorithms. Specifically, for infectious diseases, the RW algorithm works well for homogeneous contact networks, where individuals have a similar number of contacts. However, real-world contact networks are often heterogeneous (Nielsen et al., 2020), that is, the individuals vary in their number of contacts. This heterogeneity can result in size-biased samples when using the RW algorithm, i.e., over-representation of highly connected individuals (high-degree nodes) (Stein et al., 2014, Li et al., 2015).
In (Gjoka et al., 2010), the RW and MHRW algorithms are compared for sampling online social networks, specifically Facebook. The RW algorithm produces a size-biased sample, over-representing high-degree nodes, whereas the MHRW algorithm generates a sample with a degree distribution closely matching that of the underlying Facebook network. Similar observations are reported in (Leskovec and Faloutsos, 2006b) for various network types, including citation networks from the e-print arXiv repository (Leskovec and Faloutsos, 2006b), autonomous system networks (University of Oregon, ) (Internet router networks), affiliation networks from the astrophysics category on arXiv (Leskovec and Faloutsos, 2006b), and trust networks from Epinions.com (Richardson et al., 2003). We note, however, that similar comparative analyses of RW and MHRW in the context of epidemiological research have not been undertaken, to the best of our knowledge. Here, we address this gap by exploring the RW and MHRW sampling methods for infectious disease modelling.
We examine how network structures and sampling algorithms affect estimates of disease metrics. We assume a stochastic SIR model with varying transmission and recovery rates, for three network types (ER, SW and SF). Specifically, in Section 2, we quantify the size bias in the measurement of disease metrics introduced by the RW and MHRW algorithms, and in Section 3 we compare key metrics such as epidemic size, the number of secondary infections (effective reproduction number), and the time-to-infection. Finally, in Section 4, to illustrate the practical applications of our study, we analyse cattle movement network data provided by the British Cattle Movement Service (BCMS), which has hybrid properties linked to the ER, SW and SF networks. We provide a summary and directions for future research in Section 5.
@@ 3.2 Disease metrics
In the previous section, we compared the degree distribution of the samples generated using the RW and MHRW algorithms; we observed that the degree distribution of MHRW samples is closer to the UN than RW samples, due to bias towards high-degree nodes. Now, we compare the sampling algorithms in estimating three disease metrics: 1) proportion of infected nodes, 2) average number of secondary infections, and 3) time-to-infection. We simulate the SIR model on networks to estimate the disease metrics in the samples and the UN. Here, we present the results for $\gamma=1$; see C.2 for $\gamma\in\{1/7,1/14\}$.
@@ 3.2.2 Average secondary infections
Results for the time-to-infection disease metric are described in C.1. We observed that in Figure 8, estimates from both sampling algorithms (with or without duplicates) are close to the UN. Although the RW samples slightly underestimate the time-to-infection, the MHRW samples align well with the UN.
@@
The RW samples overestimated the disease metrics regardless of whether duplicate nodes were removed or retained. The size bias in degree distribution (Section 3.1), for the RW samples, leads to overestimating the number of secondary infections for all networks, as the nodes with high degrees are more likely to become infected.
@@
From an application perspective, the RW sampling algorithm provides conservative estimates of disease metrics, which is beneficial for devising interventions in cases of infectious diseases with high fatality rates. For such diseases, e.g. COVID-19 and Ebola, it is preferable to contain transmission as early as possible; hence, overestimating the number of infected individuals and the effective reproduction number (i.e., the number of secondary infections) aids the implementation of timely prevention measures. However, for diseases with low fatality rates, policymakers have more time to develop policies aligned with available resources. Here, the MHRW sampling algorithm is preferable, as its estimates are closer to the UN structure. In this context, precise estimates of disease metrics are integral for the efficient use of time and resources to contain the disease spread.
@@ 4 Cattle Movement Network
In Section 3.2, we compared the network properties and disease metrics estimated from the RW and the MHRW sampling algorithms, on three theoretical network types. Now, we implement both sampling algorithms on movement network data. We use the cattle movement data provided by the British Cattle Movement Service (BCMS), which contains information about the birth, death, and movement of the cattle.
@@ 4.1 Data Description
The data contains movement details of $1,135,502$ cattle; there are two movement records on average, per cattle. There are $46,512$ holding places, which are the nodes of the network. A static and undirected network is created with $46,512$ nodes and $159,036$ edges. Note that an edge between a pair of nodes is counted only once, in case of multiple movement records. A part of the movement network is shown in Figure 6(a); where we see a mix of large connected components, small clusters of nodes and isolated node pairs.
@@
(e) time-to-infection (days)
@@
(h) time-to-infection (days)
@@ 4.2 Results
In Figure 7, we compare three disease metrics in RW and MHRW samples relative to the cattle network, as the transmission rate $\beta$ increases, with a recovery rate $\gamma=1$, for the cattle movement network. The three disease metrics are: 1) proportion of infected nodes, 2) number of secondary infections, and 3) time-to-infection. Panel 1 (Figure 6(c), 6(d), 6(e)) shows the results for samples with duplicate nodes, while Panel 2 (Figure 6(f), 6(g), 6(h)) shows those without duplicates. The closer the disease metric estimates made from the samples are to the solid horizontal black line, the closer they are to the underlying network.
@@
In Panel 1 of Figure 7, the estimate of time-to-infection is overestimated by the MHRW samples and underestimated by the RW samples. The variability in estimates is extremely high for the MHRW samples compared to the thin band for the RW sample estimates (Figure 6(e)). Interestingly, after removing the duplicate nodes, both sampling algorithms underestimate the time to get the infection, which deviates further from the cattle network estimate as the $\beta$ value increases.
Figure 7 indicates that the estimates from MHRW samples are closer to the cattle network than the RW samples. However, the variability is significantly higher for the MHRW samples for the number of secondary infections and the time-to-infection.
@@ 5 Conclusions
In this study, we assessed the performance of the Random Walk (RW) and Metropolis-Hastings Random Walk (MHRW) sampling algorithms for estimating disease metrics across three network types: Erdős-Rényi (ER), Small-World (SW), and Scale-Free (SF). These network types were chosen because they capture essential properties of real-world networks. For instance, sexual contact networks often resemble ER networks (Holme, 2013), while the SARS-CoV-2 transmission network in Houston, Texas, has been shown to align more closely with SF networks (Fujimoto et al., 2023). Using a stochastic SIR model, we examined how size bias, caused by the over-representation of highly connected individuals, in samples generated using the RW algorithm, leads to biased estimates of disease spread. Additionally, we investigated how the inclusion of duplicate nodes influences the estimation of key disease metrics. Our analysis focused on three key disease metrics: the proportion of infected individuals, the average number of secondary infections, and the time-to-infection.
@@
Our findings (Section 3.2) highlight the suitability of the MHRW algorithm for ER and SW networks. By reducing size bias, MHRW samples consistently produced estimates that closely aligned with the underlying network’s estimates, particularly when duplicate sampled nodes were retained. In contrast, the RW algorithm produced estimates with significant overestimation, approximately $25\%$, for the proportion of infected individuals and secondary infections. Hence, the RW algorithm is less reliable for precise estimates but can be valuable for early interventions in high-fatality diseases like COVID-19 and Ebola. Conversely, the MHRW algorithm is more appropriate for diseases, like seasonal influenza. Notably, estimates of the time-to-infection metric align closely with the underlying network for both sampling algorithms.
@@
In SF networks, the inherent heterogeneity and presence of high-degree hubs introduced substantial variability in the estimates derived from both the RW and MHRW samples. The high variability in the estimated disease metrics suggests that neither method is optimal for SF networks. Instead, alternative approaches, such as separate sampling of core and peripheral nodes using algorithms like the Sampling for Large-Scale networks at low Sampling Rates (SLSR) (Jiao, 2024), may be required for more representative sampling.
@@
In Section 4, we analyzed a cattle movement network that exhibits characteristics of ER, SW, and SF networks, this further supported the findings above. MHRW sample estimates were more representative of the underlying network when duplicate nodes were retained, while RW samples significantly overestimated the proportion of the infected farms by approximately $100\%$ and the number of secondary infections by over $900\%$, reflecting size bias. Also, RW samples underestimated the time-to-infection by approximately $40\%$. These findings suggest that RW’s conservative estimates are advantageous for rapid containment measures in highly contagious diseases, such as Foot-and-Mouth Disease. Conversely, the more precise estimates from MHRW sampling are better suited for resource-efficient intervention strategies in less severe diseases, such as Bovine Viral Diarrhea.
@@
Our analysis revealed that the sensitivity of both sampling algorithms to recovery rates significantly impacts the accuracy of disease metric estimates (C.2). While the difference in the proportion of infected nodes estimated by the two sampling algorithms reduces with the recovery rate, variability in the estimates of secondary infections and time-to-infection increases with higher transmission rates.
@@
In future works, to further understand the subtle differences between the RW and MHRW sampling algorithms with respect to network structure, it would be valuable to explore the impact of selecting a specific starting node, rather than a random node, particularly for SF networks. As observed in Figure 9, the difference in disease estimates between the RW and MHRW samples decreases with the recovery rate. It would be interesting to determine the value of the recovery rate at which this change starts, as it could guide the choice of sampling algorithm. Additionally, exploring the sensitivity of these estimates to variations in network parameters, such as network size, average degree, and the power-law exponent for SF networks, would provide further insight.
@@
Conceptualization, Formal analysis, Investigation, Methodology, Validation, Visualization, Writing – original draft & editing. Katerina Kaouri: Conceptualization, Formal analysis, Investigation, Methodology, Project administration, Supervision, Validation, Writing – review & editing. Thomas E. Woolley: Conceptualization, Formal analysis, Investigation, Methodology, Project administration, Supervision, Validation, Writing – review & editing.